20n hours v. n^2 microseconds:

which has a higher order of growth? Which one is better
Answer #1

n^2 will have the higher growth and 20n is the better as it takes less time

Sample if

n=1

20 * 1=20 and 1^2=1

n=100

20 * 100= 2000 and 10^2 = 100

n=1000

20 * 1000=20000 and 1000^2=10,00,1000

so the growth n^2 will be higher compare to 20 * n

Note : Please comment below if you have concerns. I am here to help you

If you like my answer please rate and help me it is very Imp for me